OpenStudy (anonymous):

What is the solution of the system? Use substitution. y = x + 6 y = 2x A. (6, 12) B. (–12, –6) C. (–6, –12) D. (2, 4)

OpenStudy (anonymous):

A(6;12)

myininaya (myininaya):

@Palmo4ka Please provide the steps in getting the solution. It is against the code of conduct just to give answers only. So we have y=x+6 y=2x This tells us we can replace y in the first equation with 2x giving us 2x=x+6 Now subtract x on both sides

OpenStudy (anonymous):

I already gave the answer to this question http://openstudy.com/study#/updates/4f9308bce4b000ae9ec9cf8a

myininaya (myininaya):

@ericbob please don't post duplicate questions. If you aren't getting enough attention in your question, try posting a link to it in the chats. @palmo4ka and again sorry.